OpenStudy (anonymous):

A line passes through the point (-3, -3) and has a slope of 1/2. What is the equation of the line?

hartnn (hartnn):

use the slope point form: \(\Large y-y_1 = m (x-x_1)\) here m = 1/2, x1 = -3 y1 = -3

hartnn (hartnn):

just plug the values in!

OpenStudy (anonymous):

Your slope would be 1/2 and the y intercept would be -3/2

OpenStudy (anonymous):

Thank you.

hartnn (hartnn):

correct :) welcome ^_^